how do i solve this problem 3/5=(?)%times5/8
Answer by Theo(13342)

3/5 is what percent times 5/8?

let x = what percent times.

you get (3/5) = x * (5/8)

divide both sides of this equation by (5/8) to get:

x = (3/5) / (5/8)

this is equivalent to:

x = (3/5) * (8/5)

this is equivalent to:

x = (3*8) / (5*5)

this is equivalent to:

x = 24/25

24/25 is equivalent to .96

.96 is equivalent to 96%

the answer to your question is that (3/5) = 96% of (5/8)

To confirm, just translate everything to decimals.

You get:

.6 = 96% of .625

96% of .625 is equivalent to .96 * .625 is equal to .6

you are left with .6 = .6 which is true, confirming that 96% is a good figure.